OpenStudy (anonymous):

find an equation for the line tangent to y= 4-6x^2 at (5,-146)

OpenStudy (apoorvk):

differentiate to find dy/dx which gives value of slope 'm' at any particular(x",y"), which is given. now the tangent also passes through (x",y"). so write an equation for the line with slope m passing through (x",y") as: (y-y")=m(x-x")

OpenStudy (anonymous):

do you have the answer ?

OpenStudy (apoorvk):

what did you get? post and verify...

OpenStudy (apoorvk):

yeah i have my answer.. now you say what you got??

OpenStudy (anonymous):

i got Y=90x+103

OpenStudy (anonymous):

right

OpenStudy (apoorvk):

nope.. it should be y+60x=154

OpenStudy (apoorvk):

dont worry buddy, keep trying!

OpenStudy (anonymous):

no its not like what you thing its actually y= -90x+228

OpenStudy (apoorvk):

you sure?? lemme recheck